OpenStudy (anonymous):

\[\sum{x^{n}}+\sum(x/5)^{n} \Rightarrow \sum(1+1/5)x^{n} \]

OpenStudy (anonymous):

the right hand side should be \[\sum_{}^{} (1 + 1/5^{n}) * x ^{n}\]

OpenStudy (anonymous):

\[\sum{x^{n}}+\sum(x^{n}/5^{n}) \neq\sum(1+1/5)x^{n} \]

OpenStudy (anonymous):

the left hand side says (x/5) whole to the power n.